OpenStudy (anonymous):

Solve equation: 1/2lg(3x+1)-1=lg(x-3)-lg(5)

OpenStudy (anonymous):

Rules so that you can do it. \[a \times \log(b) = \log (b^{a})\] \[\log(b) - a = \log (b) - a\] \[\log(a) - \log(b) = \log(\frac{a}{b})\] \[10^{\log_{10}(a)} = a\] you will need all these equations in turn. try and get as few log terms as possible, then raise the base of the log to the log to get the bracket term

OpenStudy (anonymous):

thanks